Assume that in January 2012, the average house price in a particular area was $266444. In January 2007, the average price was $239254.

What was the annual percentage increase in selling price? (round 2 decimal places)
Answer #1
annual percentage increase in selling price = 2.18%
Dear Student
Thank you for using Chegg
Please find below the answer
Statementshowing Computations
Paticulars
Time in years = 2012 - 2007 5 Years
266,444 = 239,254 (1+r/100)^5
1.1136= (1+r/100)^5
1.0218 = 1+ r/100
.0218 =r/100
r = 2.18%